OpenStudy (help!!!!):

Find the constants A and B that makes the given equation true. How do I solve for this?

OpenStudy (help!!!!):

\[\frac{3x+15}{x^2-4x-5}=\frac{a}{x-50}+\frac{b}{x+1}\]

OpenStudy (help!!!!):

@zepdrix @dan815

OpenStudy (help!!!!):

I don't know where to start with this equation..

OpenStudy (help!!!!):

@jim_thompson5910

OpenStudy (victoriasushchik):

Multiply to get the denominators equal first. You are on the right track. 2x-9/(x^2-x-6) = A/(x-3) + B/(x+2) Multiply A by (x+2)/(x+2) and B by (x-3)/(x-3). Now all have the same denominator, so you can ignore the denominator when solving for A and B. Now you have... 2x-9 = A(x+2) + B(x-3) Distribute... 2x-9 = Ax + 2A + Bx - 3B Separate into two equations, one for X and one for your constants. 2x = Ax + Bx -9 = 2A - 3B Solve this system of equations. You can drop the x's out of the first equation because there is one in every term. 2 = A + B -9 = 2A - 3B Multiply the first equation by 3 and add to the second to cancel out the B's. 6 = 3A + 3B Add -9 = 2A - 3B + ___________ -3 = 5A A = -3/5. Plug this back into any of the two equations. 2 = A + B ===> 2 = -3/5 + B 13/5 = B So, A = -3/5. and B = 13/5.

OpenStudy (victoriasushchik):

Is This Helpful??

OpenStudy (help!!!!):

wOw thank you so much! I was taking my time to process the stuff. U r a lifesaver :D

OpenStudy (victoriasushchik):

your welcome!!!

OpenStudy (help!!!!):

I really appreciate your time and effort to help me!

OpenStudy (help!!!!):

yo @jim_thompson5910 thanks for coming by

OpenStudy (help!!!!):

Also I mistyped the problem but I understand how to solve

OpenStudy (help!!!!):

it suppose to be (x-5) not (x-50)

jimthompson5910 (jim_thompson5910):

Here is another route 2x-9 = A(x+2) + B(x-3) 2(-2)-9 = A(-2+2) + B(-2-3) .... plug in x = -2 -4-9 = A*0+B(-5) -13 = -5B -13/(-5) = B B = 13/5 Now use this to find A. You'll find that A = -3/5

OpenStudy (help!!!!):

Oh okey so just use x to get ride of one of the variables. Gotcha!

jimthompson5910 (jim_thompson5910):

@Victoriasushchik has the more effective method though